I need to wirte an expression
a person joins a book club and the first six books are free but then each costs 8.98
how much did he pay for eight books
if a bill came through with the shipment of 53.88 how would you create aqnd solve the equation
thanks

Answer by stanbon(75887)

I need to wirte an expression
a person joins a book club and the first six books are free but then each costs 8.98
how much did he pay for eight books?
Ans: 2*8.98 = $17.96
------------------------------------
if a bill came through with the shipment of 53.88 how would you create and solve the equation
---
Cost = 0*6 + 8.98(x-6), where x is the number of books bought.
53.88 = 8.98(x-6)
(x-6) = 53.88/8.98
(x-6) = 6
x = 12 books
=====================
Cheers,
Stan H.
=====================